Anna Green thinks that the recent Federal Reserve policy is going to push the interest rates up. She is considering keeping only one of the three bonds in her portfolio. She knows that bond A has a duration of 5.234, bond B has a duration of 4.867, and bond C has the following characteristics:

Par Value        1,000

Life                 5 years

Coupon Rate 5%

Discount rate   14 percent

Which one of the three bonds should she keep? Why? Explain.
Answer #1

We have to calculate duration of Bond C as follows:

Coupon rate=1,000*5%=$50

1 2 3 4=2*3 5=2*4
Year Cash Flow DF @ 14% PV Weighted average
1 50 0.8772 43.8596 43.8596
2 50 0.7695 38.4734 76.9468
3 50 0.6750 33.7486 101.2457
4 50 0.5921 29.6040 118.4161
5 1050 0.5194 545.3371 2726.6855
691.0227 3067.1537

Duration=3067.1537/691.0227

=4.4386

Anna Green thinks interest rate may rise, which means bond price will fall if Federal Reserve Policy pushes the interest rate up. Which means that she should keep a bond having lowest duration so that the price fall is lowest for increase in interest rate.

Accordingly she should keep Bond C which has the lowest duration amonst the three bonds.